# CLEAR: Calibrated Learning for Epistemic and Aleatoric Risk

### v1

@@ Appendix E Full results on real-world datasets
This appendix presents the detailed quantitative results from the benchmark experiments conducted on the 17 real-world datasets, complementing the summary findings in the main paper (Section 4.2). Description of the datasets can be found in Table 5. As a reminder, we evaluate three distinct configurations (variants) of our proposed CLEAR method alongside the baseline approaches (PCS, ALEATORIC, ALEATORIC-R). The differences among these 3 variants are considered base models. For each variant, CLEAR is applied on the already trained PCS and uncalibrated ALEATORIC-R. To recap these variants:
@@
  • Variant (c): Uses a standard PCS approach with mean-based regression models (e.g., Random Forest, XGBoost) for the epistemic component (via emprical quantiles over the estimated medians) and median prediction $\hat{f}$ (as emprical median over the estimated means), selecting the top model ($k=1$). The aleatoric component is derived from a bootstrapped ALEATORIC-R model using QRF.
@@
The subsequent subsections present the comprehensive results for each variant (a, b, c). All experiments were run across 10 different random seeds [^2] Due to a bug in pygam’s Expectile GAM [60], for the naval_propulsion data in variant (c), one of the runs failed; hence, there are 9 runs available instead of 10., and the results on the test set are presented. The subsequent tables report the average performance metric across these 10 seeds, along with the standard deviation (±$\pm\sigma$), to indicate the variability associated with data splitting. For each variant, we provide tables detailing the metrics presented in Section C.3 for all methods across all datasets at a 95% nominal coverage level. Lower values are preferable for all metrics except PICP, which should ideally be close to the target of 0.95. Additionally, summary plots (Figures 3, 7 and 8) visualize the relative NCIW and Quantile Loss performance normalized against the respective CLEAR variant baseline. Additionally, only for variant (c), we provide Table 21 with the values of $\lambda$ and $\gamma_{1}$ to provide some insights into the aleatoric/epistemic allocations of CLEAR. Finally, Figure 9 (Section E.4) provides a direct comparison of the NCIW and Quantile Loss between the three CLEAR variants themselves, using variant (a) as the reference baseline.
@@
Section E.4 shows that the primary configuration of CLEAR presented in the main text, variant (a), leverages a diverse set of quantile models for robust epistemic uncertainty estimation and generally delivers the most consistent and superior performance. As detailed in Figure 9 in the appendix, CLEAR in variant (a) tends to marginally outperform the model in variant (b), which is restricted to QXGB, and variant (c), which utilizes standard mean-based PCS models. For instance, in variant (a), QXGB was selected in approximately 64% of cases, QRF in 24%, and ExpectileGAM in the remainder of cases. In variant (c), XGBoost was chosen about 70% of the time, with other models (excluding Ridge) sharing the rest. The importance of this dynamic model selection is evident in datasets like naval_propulsion, where variants (b) and (c) can struggle due to their fixed model choices (QXGB and QRF for the aleatoric part, respectively), whereas variant (a) can adapt by selecting, for example, ExpectileGAM. This highlights the stability advantage of CLEAR (a) and the benefit of PCS’s dynamic model choice, especially when the aleatoric component can also adapt. Nevertheless, the CLEAR framework overall demonstrates considerable robustness even with these alternative model choices. For instance, variant (b) still provides NCIW reductions of approximately 17.3% and 4.9% over its corresponding PCS and ALEATORIC-R baselines, respectively (Figure 7). A similar advantage is observed for variant (c) (Figure 8), which reduces NCIW by about 7.6% against ALEATORIC-R, the best-performing variant of CQR that uses our novel residual-based technique.
@@
Overall, variant (a) archives the best results (see Figure 9). Therefore, we recommend to use CLEAR variant (a). The experiments in variant (a) are fair, since each of the competing methods uses the same base model, which is selected per data set based on the RMSE on the validation data set. See Section E.1 for the results.
@@
Variant (b) is the simplest to understand, easiest to implement, and computationally cheapest variant and provides maximal fairness, making it particularly scientifically sound. Each method simply uses QXGB as base model without any model selection step. Strictly speaking, variant (b) is the only variant where ALEATORIC and ALEATORIC-R are fully conformal, since variant (b) uses the calibration data set only for calibration, while variant (a) and (c) reuse the validation set used for model selection as calibration data set. However, in practice, we observe this re-usage does not hurt calibration in any significant way (Table 6 and Table 16). See Section E.2 for the results of variant (b).
@@
Variant (c) uses the same set of base models as suggested by the authors of PCS uncertainty quantification [23]. [23] conducted extensive experiments showing that PCS uncertainty quantification substantially outperforms popular conformal baselines such as split conformal regression [101], Studentized conformal regression [101], and Majority Vote [102] (by more than 20% on average in terms of interval width). Our experiments showing that CLEAR (a) outperforms CLEAR (c) (Figure 9) and CLEAR (c) outperforms PCS uncertainty quantification (c) (Section E.3), together with the experiments by [23], strongly suggests that CLEAR clearly outperforms these popular conformal baselines. See Section E.3 for the results of variant (c).
@@ E.1 Variant (a)
Table 7: Variant (a) NIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
naval_propulsion 6.1e-04 $\pm$ 6.5e-06 6.2e-04 $\pm$ 6.8e-06 6.0e-04 $\pm$ 6.7e-06 7.2e-04 $\pm$ 1.4e-05 6.0e-04 $\pm$ 7.3e-06 6.2e-04 $\pm$ 5.8e-06 6.1e-04 $\pm$ 5.8e-06
parkinsons 0.227 $\pm$ 0.010 0.321 $\pm$ 0.013 0.250 $\pm$ 0.014 0.312 $\pm$ 0.021 0.303 $\pm$ 0.022 0.228 $\pm$ 0.008 0.225 $\pm$ 0.009
@@
sulfur 0.109 $\pm$ 0.011 0.112 $\pm$ 0.010 0.107 $\pm$ 0.010 0.131 $\pm$ 0.017 0.122 $\pm$ 0.015 0.108 $\pm$ 0.011 0.108 $\pm$ 0.011
superconductor 0.197 $\pm$ 0.021 0.233 $\pm$ 0.025 0.197 $\pm$ 0.023 0.248 $\pm$ 0.028 0.329 $\pm$ 0.042 0.195 $\pm$ 0.021 0.194 $\pm$ 0.021
Table 8: Variant (a) QUANTILELOSS at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
naval_propulsion 1.5e-05 $\pm$ 2.2e-07 1.5e-05 $\pm$ 2.7e-07 1.5e-05 $\pm$ 2.8e-07 1.7e-05 $\pm$ 2.8e-07 1.6e-05 $\pm$ 3.5e-07 1.5e-05 $\pm$ 2.2e-07 1.5e-05 $\pm$ 2.4e-07
parkinsons 0.178 $\pm$ 0.009 0.204 $\pm$ 0.008 0.215 $\pm$ 0.012 0.224 $\pm$ 0.012 0.251 $\pm$ 0.014 0.185 $\pm$ 0.010 0.178 $\pm$ 0.010
@@
sulfur 2.0e-03 $\pm$ 1.3e-04 2.1e-03 $\pm$ 1.2e-04 2.3e-03 $\pm$ 1.1e-04 2.3e-03 $\pm$ 1.9e-04 3.5e-03 $\pm$ 2.0e-04 2.0e-03 $\pm$ 1.3e-04 2.0e-03 $\pm$ 1.3e-04
superconductor 0.490 $\pm$ 0.018 0.512 $\pm$ 0.015 0.539 $\pm$ 0.024 0.608 $\pm$ 0.023 0.861 $\pm$ 0.036 0.490 $\pm$ 0.018 0.491 $\pm$ 0.018
Table 9: Variant (a) NCIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
naval_propulsion 6.1e-04 $\pm$ 8.4e-06 6.1e-04 $\pm$ 7.8e-06 6.0e-04 $\pm$ 5.6e-06 7.1e-04 $\pm$ 1.5e-05 6.0e-04 $\pm$ 5.6e-06 6.1e-04 $\pm$ 9.7e-06 6.1e-04 $\pm$ 6.7e-06
parkinsons 0.226 $\pm$ 0.010 0.321 $\pm$ 0.013 0.248 $\pm$ 0.008 0.316 $\pm$ 0.032 0.302 $\pm$ 0.013 0.228 $\pm$ 0.007 0.225 $\pm$ 0.009
@@
sulfur 0.109 $\pm$ 0.010 0.111 $\pm$ 0.009 0.107 $\pm$ 0.009 0.128 $\pm$ 0.013 0.129 $\pm$ 0.011 0.108 $\pm$ 0.011 0.107 $\pm$ 0.011
superconductor 0.195 $\pm$ 0.021 0.232 $\pm$ 0.025 0.194 $\pm$ 0.021 0.247 $\pm$ 0.028 0.314 $\pm$ 0.035 0.194 $\pm$ 0.021 0.194 $\pm$ 0.021
Table 10: Variant (a) INTERVALSCORELOSS at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
naval_propulsion 1.2e-03 $\pm$ 1.8e-05 1.2e-03 $\pm$ 2.2e-05 1.2e-03 $\pm$ 2.2e-05 1.4e-03 $\pm$ 2.3e-05 1.3e-03 $\pm$ 2.8e-05 1.2e-03 $\pm$ 1.8e-05 1.2e-03 $\pm$ 1.9e-05
parkinsons 14.235 $\pm$ 0.752 16.303 $\pm$ 0.641 17.210 $\pm$ 0.950 17.947 $\pm$ 0.957 20.088 $\pm$ 1.130 14.779 $\pm$ 0.840 14.274 $\pm$ 0.762
@@
sulfur 0.161 $\pm$ 0.011 0.172 $\pm$ 0.010 0.187 $\pm$ 0.008 0.186 $\pm$ 0.015 0.280 $\pm$ 0.016 0.161 $\pm$ 0.011 0.161 $\pm$ 0.011
superconductor 39.230 $\pm$ 1.437 40.925 $\pm$ 1.177 43.103 $\pm$ 1.911 48.630 $\pm$ 1.867 68.891 $\pm$ 2.882 39.196 $\pm$ 1.455 39.314 $\pm$ 1.475
E.2 Variant (b)
Figure 7: Quantile loss and NCIW performance of different methods (CLEAR, PCS, ALEATORIC, ALEATORIC-R) for variant (b), over 10 seeds normalized relative to CLEAR (baseline = 1.0). Lower values indicate better performance. The inset boxplot shows the % improvement relative to the CLEAR baseline ±$\pm 1\sigma$. Values inside each subplot represent the mean improvement across all datasets.
@@ E.2 Variant (b)
Table 12: Variant (b) NIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
superconductor 0.218 $\pm$ 0.023 0.309 $\pm$ 0.033 0.229 $\pm$ 0.025 0.240 $\pm$ 0.026 0.354 $\pm$ 0.043 0.223 $\pm$ 0.024 0.215 $\pm$ 0.024
Table 13: Variant (b) QUANTILELOSS at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
superconductor 0.522 $\pm$ 0.018 0.647 $\pm$ 0.019 0.573 $\pm$ 0.023 0.607 $\pm$ 0.025 0.891 $\pm$ 0.035 0.522 $\pm$ 0.018 0.522 $\pm$ 0.019
Table 14: Variant (b) NCIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
superconductor 0.217 $\pm$ 0.023 0.308 $\pm$ 0.032 0.226 $\pm$ 0.023 0.238 $\pm$ 0.026 0.347 $\pm$ 0.037 0.222 $\pm$ 0.023 0.215 $\pm$ 0.023
Table 15: Variant (b) INTERVALSCORELOSS at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
superconductor 41.786 $\pm$ 1.441 51.742 $\pm$ 1.524 45.854 $\pm$ 1.855 48.599 $\pm$ 2.033 71.287 $\pm$ 2.788 41.754 $\pm$ 1.448 41.788 $\pm$ 1.549
E.3 Variant (c)
Figure 8: Quantile loss and NCIW performance of different methods (CLEAR, PCS, ALEATORIC, ALEATORIC-R) for variant (c), over 10 seeds normalized relative to CLEAR (baseline = 1.0). Lower values indicate better performance. The inset boxplot shows the % improvement relative to the CLEAR baseline ±$\pm 1\sigma$. Values inside each subplot represent the mean improvement across all datasets.

### v2

@@ Appendix F CLEAR with PCS and CQR: Results on Real-World Data
This appendix presents the detailed quantitative results from the benchmark experiments conducted on the 17 real-world datasets, complementing the summary findings in the main paper (Section 4.2). As a reminder, we evaluate three distinct configurations (variants) of our proposed CLEAR method alongside the baseline approaches (PCS, ALEATORIC, ALEATORIC-R). The differences among these 3 variants are considered base models. For each variant, CLEAR is applied on the already trained PCS and uncalibrated ALEATORIC-R. To recap these variants:
@@
  • Variant (c): Uses a standard PCS approach with mean-based regression models (e.g., Random Forest, XGBoost) for the epistemic component (via empirical quantiles over the estimated medians) and median prediction $\hat{f}$ (as empirical median over the estimated means), selecting the top model ($k=1$). The aleatoric component is derived from a bootstrapped ALEATORIC-R model using QRF.
@@
The subsequent subsections present the comprehensive results for each variant (a, b, c). All experiments were run across 10 different random seeds [^7] Due to a bug in pygam’s Expectile GAM [39], for the naval_propulsion data in variant (c), one of the runs failed; hence, there are 9 runs available instead of 10., and the results on the test set are presented. The subsequent tables report the average performance metric across these 10 seeds, along with the standard deviation ($\pm\sigma$), to indicate the variability associated with data splitting. For each variant, we provide tables detailing the metrics presented in Section D.5 for all methods across all datasets at a 95% nominal coverage level. Lower values are preferable for all metrics except PICP, which should ideally be close to the target of 0.95. Additionally, summary plots (Figures 3, 7 and 8) visualize the relative NCIW and Quantile Loss performance normalized against the respective CLEAR variant baseline. Additionally, only for variant (c), we provide Table 27 with the values of $\lambda$ and $\gamma_{1}$ to provide some insights into the aleatoric/epistemic allocations of CLEAR. Finally, Figure 9 (Section F.4) provides a direct comparison of the NCIW and Quantile Loss between the three CLEAR variants themselves, using variant (a) as the reference baseline.
@@
Section F.4 shows that the primary configuration of CLEAR presented in the main text, variant (a), leverages a diverse set of quantile models for robust epistemic uncertainty estimation and generally delivers the most consistent and superior performance. As detailed in Figure 9 in the appendix, CLEAR in variant (a) tends to marginally outperform the model in variant (b), which is restricted to QXGB, and variant (c), which utilizes standard mean-based PCS models. For instance, in variant (a), QXGB was selected in approximately 64% of cases, QRF in 24%, and ExpectileGAM in the remainder of cases. In variant (c), XGBoost was chosen about 70% of the time, with other models (excluding Ridge) sharing the rest. The importance of this dynamic model selection is evident in datasets like naval_propulsion, where variants (b) and (c) can struggle due to their fixed model choices (QXGB and QRF for the aleatoric part, respectively), whereas variant (a) can adapt by selecting, for example, ExpectileGAM. This highlights the stability advantage of CLEAR (a) and the benefit of PCS’s dynamic model choice, especially when the aleatoric component can also adapt. Nevertheless, the CLEAR framework overall demonstrates considerable robustness even with these alternative model choices. For instance, variant (b) still provides NCIW reductions of approximately 17.3% and 4.9% over its corresponding PCS and ALEATORIC-R baselines, respectively (Figure 7). A similar advantage is observed for variant (c) (Figure 8), which reduces NCIW by about 7.6% against ALEATORIC-R, the best-performing variant of CQR that uses our novel residual-based technique.
@@
Overall, variant (a) archives the best results (see Figure 9). Therefore, we recommend to use CLEAR variant (a). The experiments in variant (a) are fair, since each of the competing methods uses the same base model, which is selected per data set based on the RMSE on the validation data set. See Section F.1 for the results.
@@
Variant (b) is the simplest to understand, easiest to implement, and computationally cheapest variant and provides maximal fairness, making it particularly scientifically sound. Each method simply uses QXGB as base model without any model selection step. Strictly speaking, variant (b) is the only variant where ALEATORIC and ALEATORIC-R are fully conformal, since variant (b) uses the calibration data set only for calibration, while variant (a) and (c) reuse the validation set used for model selection as calibration data set. However, in practice, we observe this re-usage does not hurt calibration in any significant way (Table 12 and Table 22). See Section F.2 for the results of variant (b).
@@
Variant (c) uses the same set of base models as suggested by the authors of PCS uncertainty quantification as [30]. They conducted extensive experiments showing that PCS uncertainty quantification substantially outperforms popular conformal baselines such as split conformal regression [111], Studentized conformal regression [111], and Majority Vote [112] (by more than 20% on average in terms of interval width). Our experiments showing that CLEAR (a) outperforms CLEAR (c) (Figure 9) and CLEAR (c) outperforms PCS uncertainty quantification (c) (Section F.3), together with the experiments by [30], strongly suggests that CLEAR clearly outperforms these popular conformal baselines. See Section F.3 for the results of variant (c).
@@ F.1 Variant (a)
Table 13: Variant (a) NIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
naval_propulsion 6.1e-04 $\pm$ 6.3e-06 6.2e-04 $\pm$ 6.5e-06 6.0e-04 $\pm$ 6.4e-06 7.2e-04 $\pm$ 1.3e-05 6.0e-04 $\pm$ 7.0e-06 6.2e-04 $\pm$ 5.7e-06 6.1e-04 $\pm$ 5.6e-06
parkinsons 0.227 $\pm$ 0.010 0.321 $\pm$ 0.013 0.249 $\pm$ 0.014 0.312 $\pm$ 0.021 0.303 $\pm$ 0.022 0.228 $\pm$ 0.008 0.225 $\pm$ 0.008
@@
sulfur 0.109 $\pm$ 0.010 0.112 $\pm$ 0.010 0.107 $\pm$ 0.010 0.131 $\pm$ 0.017 0.122 $\pm$ 0.015 0.108 $\pm$ 0.011 0.108 $\pm$ 0.011
superconductor 0.196 $\pm$ 0.022 0.233 $\pm$ 0.025 0.196 $\pm$ 0.023 0.248 $\pm$ 0.028 0.329 $\pm$ 0.042 0.194 $\pm$ 0.022 0.194 $\pm$ 0.021
Table 14: Variant (a) Quantile Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
naval_propulsion 1.5e-05 $\pm$ 2.1e-07 1.5e-05 $\pm$ 2.6e-07 1.5e-05 $\pm$ 2.6e-07 1.7e-05 $\pm$ 2.7e-07 1.6e-05 $\pm$ 3.5e-07 1.5e-05 $\pm$ 2.1e-07 1.5e-05 $\pm$ 2.3e-07
parkinsons 0.178 $\pm$ 0.010 0.202 $\pm$ 0.007 0.215 $\pm$ 0.012 0.224 $\pm$ 0.012 0.251 $\pm$ 0.014 0.185 $\pm$ 0.011 0.178 $\pm$ 0.010
@@
sulfur 2.0e-03 $\pm$ 1.4e-04 2.1e-03 $\pm$ 1.3e-04 2.3e-03 $\pm$ 1.1e-04 2.3e-03 $\pm$ 1.9e-04 3.5e-03 $\pm$ 2.0e-04 2.0e-03 $\pm$ 1.4e-04 2.0e-03 $\pm$ 1.4e-04
superconductor 0.490 $\pm$ 0.018 0.511 $\pm$ 0.014 0.538 $\pm$ 0.023 0.608 $\pm$ 0.023 0.861 $\pm$ 0.036 0.490 $\pm$ 0.018 0.491 $\pm$ 0.018
Table 15: Variant (a) NCIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
naval_propulsion 6.1e-04 $\pm$ 7.9e-06 6.1e-04 $\pm$ 7.4e-06 6.0e-04 $\pm$ 5.5e-06 7.1e-04 $\pm$ 1.5e-05 5.9e-04 $\pm$ 6.1e-06 6.1e-04 $\pm$ 9.3e-06 6.1e-04 $\pm$ 6.4e-06
parkinsons 0.225 $\pm$ 0.012 0.321 $\pm$ 0.013 0.247 $\pm$ 0.007 0.316 $\pm$ 0.032 0.302 $\pm$ 0.013 0.227 $\pm$ 0.007 0.224 $\pm$ 0.008
@@
sulfur 0.109 $\pm$ 0.010 0.111 $\pm$ 0.008 0.106 $\pm$ 0.009 0.128 $\pm$ 0.013 0.129 $\pm$ 0.011 0.108 $\pm$ 0.011 0.107 $\pm$ 0.010
superconductor 0.195 $\pm$ 0.021 0.231 $\pm$ 0.025 0.194 $\pm$ 0.021 0.247 $\pm$ 0.028 0.314 $\pm$ 0.035 0.194 $\pm$ 0.021 0.194 $\pm$ 0.021
Table 16: Variant (a) Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
naval_propulsion 1.2e-03 $\pm$ 1.7e-05 1.2e-03 $\pm$ 2.1e-05 1.2e-03 $\pm$ 2.1e-05 1.4e-03 $\pm$ 2.2e-05 1.3e-03 $\pm$ 2.8e-05 1.2e-03 $\pm$ 1.7e-05 1.2e-03 $\pm$ 1.8e-05
parkinsons 14.221 $\pm$ 0.792 16.142 $\pm$ 0.582 17.188 $\pm$ 0.985 17.947 $\pm$ 0.957 20.088 $\pm$ 1.130 14.762 $\pm$ 0.882 14.259 $\pm$ 0.786
@@
sulfur 0.161 $\pm$ 0.011 0.172 $\pm$ 0.011 0.187 $\pm$ 0.008 0.186 $\pm$ 0.015 0.280 $\pm$ 0.016 0.161 $\pm$ 0.011 0.161 $\pm$ 0.011
superconductor 39.217 $\pm$ 1.419 40.851 $\pm$ 1.157 43.049 $\pm$ 1.848 48.630 $\pm$ 1.867 68.891 $\pm$ 2.882 39.182 $\pm$ 1.438 39.296 $\pm$ 1.444
F.2 Variant (b)
Figure 7: Quantile loss and NCIW performance of different methods (CLEAR, PCS, ALEATORIC, ALEATORIC-R) for variant (b), over 10 seeds normalized relative to CLEAR (baseline = 1.0). Lower values indicate better performance. The inset boxplot shows the % improvement relative to the CLEAR baseline $\pm 1\sigma$. Values inside each subplot represent the mean improvement across all datasets.
@@ F.2 Variant (b)
Table 18: Variant (b) NIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
superconductor 0.219 $\pm$ 0.023 0.309 $\pm$ 0.033 0.228 $\pm$ 0.026 0.243 $\pm$ 0.027 0.354 $\pm$ 0.042 0.223 $\pm$ 0.025 0.216 $\pm$ 0.024
Table 19: Variant (b) Quantile Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
superconductor 0.523 $\pm$ 0.018 0.648 $\pm$ 0.020 0.573 $\pm$ 0.024 0.611 $\pm$ 0.025 0.892 $\pm$ 0.035 0.522 $\pm$ 0.018 0.523 $\pm$ 0.020
Table 20: Variant (b) NCIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
superconductor 0.218 $\pm$ 0.022 0.308 $\pm$ 0.032 0.226 $\pm$ 0.022 0.241 $\pm$ 0.028 0.347 $\pm$ 0.036 0.223 $\pm$ 0.024 0.216 $\pm$ 0.024
Table 21: Variant (b) Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
@@
superconductor 41.801 $\pm$ 1.407 51.826 $\pm$ 1.568 45.860 $\pm$ 1.883 48.899 $\pm$ 2.037 71.323 $\pm$ 2.803 41.747 $\pm$ 1.460 41.821 $\pm$ 1.571
F.3 Variant (c)
Figure 8: Quantile loss and NCIW performance of different methods (CLEAR, PCS, ALEATORIC, ALEATORIC-R) for variant (c), over 10 seeds normalized relative to CLEAR (baseline = 1.0). Lower values indicate better performance. The inset boxplot shows the % improvement relative to the CLEAR baseline $\pm 1\sigma$. Values inside each subplot represent the mean improvement across all datasets.